OpenStudy (anonymous):

6 = X plus 2 over 3?

jhonyy9 (jhonyy9):

6=x+2/3 18=3x+2 16=3x x=16/3

jhonyy9 (jhonyy9):

turabq - why over 5 ?

OpenStudy (anonymous):

@jhonny9 awww. its a misspelling. i should type more carefully. thanks for the feedback.

jhonyy9 (jhonyy9):

give medal for !!! thank you

OpenStudy (anonymous):

+1 medal